# EgoDTM: Towards 3D-Aware Egocentric Video-Language Pretraining

### v1

@@ 3.3 Spatial-Aware Textual Enrichment
Previous egocentric video-language pretraining frameworks typically learn from short descriptions composed of verbs and nouns. To foster 3D-aware pretraining, we argue that captions should contain more spatial-temporal information like HOI positions, object shapes, and movements. We therefore enrich the captions with visual cues by employing a detect-track-generate pipeline tailored for egocentric videos using foundation models [56, 54, 13].
Detect. We use the detector [56], HOID, finetuned on egocentric data to detect hands, objects, and their contact states for a given frame. The detector $f(\cdot)$ receives $V_{i}$ as input and outputs diverse information $\{(b_{j},c_{j},t,s)\}_{j}$, where $b_{j}$ is the bounding boxes with confidence scores $c_{j}$, $t$ is object type, and $s$ is contact state of hand.
Track. Since predictions of HOID lack temporal consistency, we address this issue by tracking across the video. We utilize SAM2 [54] $g(\cdot)$, a foundation model for promptable segmentation and tracking. The initial tracking frame $V^{*}$ is selected when the hands make contact with objects. The HOI masks of this frame are obtained by prompting SAM2 with HOI boxes $\{m^{*}_{j}\}_{j}=g(V^{*},\{b^{*}_{j}\}_{j})$, where the bounding boxes serve as negative prompts for each other. Finally, the model tracks HOI masks throughout all frames to acquire the temporal consistent masks $M=(\{m_{1j}\}_{j},\dots,\{m_{Fj}\}_{j})$.
Generate. As suggested by previous works [34, 19], LLMs are capable of understanding visual cues like bounding boxes. We use an LLM $h(\cdot)$ to comprehend the HOI spatial information and original captions, then generate spatial-aware captions $T^{\text{aug}}=h(T,M)$.
Figure 3: Qualitative results of depth estimation from EgoDTM and DepthAnythingv2 [69](DAv2) on datasets including in-domain but unseen Ego4D validation set [20], out-of-domain and unseen data of EK100 [10], MECCANO [51], and H2O [26]. Note that DAv2 operates with a high resolution of 512p, while EgoDTM uses a lower resolution input of 224p and generates a depth map at a resolution of 56p. Despite the lower resolution input, EgoDTM demonstrates intuitive generalization across unseen egocentric datasets with diverse environments, illuminations, backgrounds, and varying HOI object sizes.
Equipped with the enriched video-text pairs, our video-text contrastive learning objective is represented as:
@@ 3.4 Training Strategy
Pretraining data. Our pretraining data consists of four million (video, text) pairs, with each video approximately 1 second long. We generate enhanced text descriptions and depth maps specifically for videos featuring hand-object interactions, resulting in two million (video, enriched texts, depths) triplets. Specifically, for each video, we extract eight depth maps using the DAv2-Large [69] model. The HOI boxes are detected by HOID [56], which is finetuned on egocentric data. The HOI masks are segmented and tracked using the same unified SAM2-Large model [54]. The final enriched text descriptions are generated using the DeepSeek-LLM-200B model [13].
@@
Method Supervision Text Param mAP (%) nDCG (%) mean-acc top1-acc Inter-Acc Intra-Acc
@@
EgoVLP [33] Text 4M 172M 26.0 20.6 23.3 28.8 27.0 27.9 - - 90.6 57.2
EgoVLPv2 [47] Text 4M 363M 35.1 26.6 30.8 33.7 30.4 32.0 30.9 35.1 91.0 60.9
LaViLa [78] Text 35M 169M 35.1 26.6 30.8 33.7 30.4 32.0 30.9 35.1 93.6 59.1
AVION [77] Text 35M 142M 37.1 28.7 32.9 34.4 31.0 32.7 38.6 42.3 94.4 62.1
HelpingHands [72] Text, Box 35M 206M 35.6 26.8 31.2 34.7 31.7 33.2 29.4 35.3 93.2 58.8
HENASY [43] Text, Box 35M 277M 35.5 27.1 31.3 34.6 31.7 33.2 29.6 35.9 94.1 61.3
EgoDTM (ours) Text, Depth 6M 151M 37.9 29.1 33.5 34.8 31.9 33.4 40.2 43.2 94.6 63.6
Table 1: Comparison with state-of-the-art methods on zero-shot video-text retrieval and action recognition. All methods use the same 4M videos from Ego4D for pretraining and have similar hidden dimension sizes. The numbers of the method with are sourced from [43].
@@ 4 Experiments
To evaluate our EgoDTM, we conduct experiments from three perspectives: short video understanding (video-text retrieval and action recognition), 3D space understanding (depth estimation and robot manipulation), and long video understanding (natural language query). These evaluations cover seven benchmarks across five datasets. In the following sections, we detail the experimental setups (Section 4.1), main results (Section 4.2), ablation studies (Section 4.3), and further analyses (Section 4.4).
@@ 4.1 Benchmarks and Settings
Zero-shot video-text retrieval (ZS-VTR). To assess video-text alignment, we conduct zero-shot retrieval evaluations on text-to-video multi-choice retrieval on Ego4D [20] (EgoMCQ) and multi-instance video-text retrieval on Epic-Kitchens-100 [10] (EK-100-MIR). Following [33, 10], we use mean Average Precision (mAP) and the normalized Discounted Cumulative Gain (nDCG) as metrics for EK-100-MIR and accuracy for EgoMCQ.
Zero-shot action recognition (ZS-AR). Action recognition is conducted in a video-to-text retrieval manner. We experiment on the EGTEA [28] dataset, requiring models to recognize 106 classes of cooking activities. The metrics include mean accuracy and top-1 accuracy across test splits.
Depth estimation (DE). To assess the model’s 3D awareness, we conduct depth estimation by finetuning our model on the H2O [26] dataset. Following [17], we add a prediction head composed of a linear layer and an upsampling convolution layer on top of the frozen video encoder. Similar to [17, 16], the metrics are threshold accuracy ($\delta_{i}$): percentage of pixels $\tilde{D}_{j}$ in ground-truth depth maps that satisfy $\text{max}(\frac{D_{{\rm gt}j}}{\tilde{D}_{j}},\frac{\tilde{D}_{j}}{D_{{\rm gt}\ j}})=\delta_{i}<thr^{i}$ for $thr=1.25,i\in\{1,2,3\}$, and RMSE: $\|\tilde{D}-{D}_{\rm gt}\|_{2}$. The scale-invariant metrics are shifted and normalized from scale-aware metrics.
@@
ConvNext [37] 0.721 0.965 0.991 0.644 0.727 0.969 0.996 0.593
CLIP [49] 0.795 0.966 0.988 0.624 0.811 0.976 0.994 0.559
EgoVLP [33] 0.778 0.954 0.989 0.610 0.853 0.977 0.996 0.497
LaViLa [78] 0.801 0.954 0.987 0.598 0.811 0.964 0.993 0.552
AVION [77] 0.786 0.960 0.991 0.606 0.812 0.969 0.996 0.543
@@
Table 2: Comparisons of depth estimation task on H2O dataset.
Robot manipulation (RM). The robot must learn to accurately perceive and interact in 3D spaces to accomplish daily tasks. Following previous works [42, 24], we evaluate the visual representations as frozen perception modules for downstream policy learning within the Franka Kitchens simulation environment [21]. Five tasks are adopted: turn knob (TK), open door (OD), flip switch (FS), open microwave (OM), and slide door (SD). We use success rate as the evaluation metric.
Natural language queries (NLQ). The NLQ task localizes time intervals in a long video given a language query. We experiment on the EgoNLQ task on Ego4D [20]. To fairly evaluate different pretrained visual representations, we extract the visual features by video encoders and text features by the same pretrained BERT [14], then train a VSLNet [73] to solve this task. The evaluation metrics are “Rn@m”, where $n\in\{1,5\}$ and $m\in\{0.3,0.5\}$, presenting the percentage of at least one of the top-$n$ predicted intervals having IoU greater than $m$.
Pretraining details. The dual-encoders are initialized by the checkpoint pretrained on the four million video-text pairs from Ego4D. EgoDTM is then trained for two epochs on 8*A800 GPUs, which requires approximately 10 hours and a learning rate of 3e-5. The depth-aware transformer decoder comprises six layers. The hidden dimension of the dual encoders is 768, while the 3D-aware decoder uses a dimension of 256 for efficient design. We use frames with 224p as input and 56p as output of the depth maps. Consequently, our 3D-aware decoder only has 9M parameters, and the batch size is set to 4096.
@@
R3M [42] 53.3% 50.7% 59.3% 86.3% 97.7% 69.4%
MPI [24] 83.3% 54% 44.5% 93.5% 100% 75%
ResNet [22] 28% 18% 26.7% 50% 75.5% 39.7%
CLIP [49] 26.3% 13% 24.7% 41.7% 86.3% 38.4%
LaViLa [78] 48% 26% 22.5% 69% 94.5% 52%
@@
4.2 Main Results
EK-100-MIR EgoMCQ
@@ 4.2 Main Results
(d) Model size. Larger model sizes bring more performance gain, especially on intra accuracy of EgoMCQ.
EK-100-MIR EgoMCQ
@@
Table 4: Zero-shot ablation studies. Models are pretrained on 4M video-text pairs from Ego4D and evaluated with zero-shot video-text retrieval on Epic-Kitchens-100-MIR [10] and EgoMCQ [33]. Unless specified, the default setting includes: ViT-B/16 backbone, batch size of 2048, random substitution of enriched texts, both $\mathcal{L}_{\rm depth}$ and $\mathcal{L}_{\rm augvtc}$ as losses, 8 depth queries, and 56p depth maps.
Zero-shot video-text retrieval. In Table 1, EgoDTM outperforms models like LaViLa [78] and AVION [77]. For example, on EK100-MIR, EgoDTM outperforms the state-of-the-art AVION model by +0.9% in mAP and +0.3% in nDCG, while also achieving a +1.5% improvement in intra accuracy on EgoMCQ. This indicates that captions generated based on visual cues from our data curation pipeline are more accurate and informative than captions generated from visual-conditioned LLMs or simply text rewriting. Besides, our model surpasses HelpingHands [72] and HENASY [43], which rely on noisy HOI detection supervision with extra parameters in the visual encoder, indicating that depth modality can potentially offer larger merits for video-language models.
Zero-shot action recognition  Table 1 illustrates the superior performance of EgoDTM on EGTEA. While previous works have demonstrated that depth modality enhances traditional action recognition tasks [74, 35], their applicability has been largely constrained to specific scenarios and closed-set datasets. Our results extend this understanding, showing that video-language models can effectively leverage depth modality to learn more generalizable video representations.
Depth estimation. We evaluate the ability of visual-language models to infer depth from images in Table 2. Our model surpasses all previous video-language models across most metrics by a large margin. In particular, our method improves the scale-aware RMSE by approximately 9.8% over the second-best result achieved by LaViLa. Interestingly, EgoVLP performs better in scale-invariant $\delta_{1}$ metric, suggesting that it latently captures finer-grained pixel-level details. We hypothesize that this advantage stems from EgoVLP’s scene-aware negative sampling strategy, which samples video-text pairs captured from the same environments into the same batch, thereby enhancing its ability to learn specific details.
Robot manipulation. As shown in Table 3, EgoDTM consistently outperforms pretrained visual-language models such as CLIP [49] and LaViLa [78] by +20.2% and +6.6%, respectively, demonstrating stronger spatial perception in visual representations. Additionally, EgoDTM performs competitively with specialized robot learning methods on certain tasks, such as “turn knob” but underperforms on others, like “open microwave”. A possible reason is that our model, pretained on depth and text, may overfit to real-worold scenarios, whereas methods like R3M [42] leverage self-supervised pretraining, which could provide better generalization to diverse manipulation tasks.
Method R1@0.3 R5@0.3 R1@0.5 R5@0.5
EgoVLP [33] 6.32 13.84 3.41 8.80
LaViLa [78] 7.12 14.82 3.87 9.55
AVION [77] 7.33 14.89 4.31 9.53
EgoDTM (ours) 8.13 16.11 4.83 10.30
Table 5: Comparisons of NLQ task on Ego4D dataset.
@@ 4.3 Ablation Study
Ablation results are shown in Table 4.
EgoDTM’s components. Combining video-text matching with depth estimation enhances each learning objective. EgoDTM leverages depth maps to capture object relations, while textual descriptions enriched with shape and movement details can benefit more from depth information.

### v2

@@ 3.3 Spatial-Aware Textual Enrichment
Previous egocentric video-language pretraining frameworks typically learn from short descriptions composed of verbs and nouns. To foster 3D-aware pretraining, we argue that captions should contain more spatial-temporal information like HOI positions, object shapes, and movements. We therefore enrich the captions with visual cues by employing a detect-track-generate pipeline tailored for egocentric videos using foundation models hoidetector; sam2; deepseekllm.
Detect. We use the detector hoidetector, HOID, finetuned on egocentric data to detect hands, objects, and their contact states for a given frame. The detector $f(\cdot)$ receives $V_{i}$ as input and outputs diverse information $\{(b_{j},c_{j},t,s)\}_{j}$, where $b_{j}$ is the bounding boxes with confidence scores $c_{j}$, $t$ is object type, and $s$ is contact state of hand.
Track. Since predictions of HOID lack temporal consistency, we address this issue by tracking across the video. We utilize SAM2 sam2 $g(\cdot)$, a foundation model for promptable segmentation and tracking. The initial tracking frame $V^{*}$ is selected when the hands make contact with objects. The HOI masks of this frame are obtained by prompting SAM2 with HOI boxes $\{m^{*}_{j}\}_{j}=g(V^{*},\{b^{*}_{j}\}_{j})$, where the bounding boxes serve as negative prompts for each other. Finally, the model tracks HOI masks throughout all frames to acquire the temporal consistent masks $M=(\{m_{1j}\}_{j},\dots,\{m_{Fj}\}_{j})$.
Generate. As suggested by previous works liu2023llava; feng2023layoutgpt, LLMs are capable of understanding visual cues like bounding boxes. We use an LLM $h(\cdot)$ to comprehend the HOI spatial information and original captions, then generate spatial-aware captions $T^{\text{aug}}=h(T,M)$.
Equipped with the enriched video-text pairs, our new video-text contrastive learning objective is:
@@ 3.4 Training Strategy
Pretraining data. Our pretraining data consists of four million (video, text) pairs, with each video approximately 1 second long. We generate enhanced text descriptions and depth maps specifically for videos featuring hand-object interactions, resulting in two million (video, enriched texts, depths) triplets. Specifically, for each video, we extract eight depth maps using the DAv2-Large depthanythingv2 model. The HOI boxes are detected by HOID hoidetector, which is finetuned on egocentric data. The HOI masks are segmented and tracked using the same unified SAM2-Large model sam2. The final enriched text descriptions are generated using the DeepSeek-LLM-200B model deepseekllm.
Table 1: Comparison with state-of-the-art methods on zero-shot video-text retrieval and action recognition. All methods use the same 4M videos from Ego4D for pretraining and have similar hidden dimension sizes. The numbers of the method with are sourced from phan2024henasy.
@@
Method mAP (%) nDCG (%) mean top1 Inter Intra
@@
EgoVLP lin2022egovlp 26.0 20.6 23.3 28.8 27.0 27.9 - - 90.6 57.2
EgoVLPv2 pramanick2023egovlpv2 35.1 26.6 30.8 33.7 30.4 32.0 30.9 35.1 91.0 60.9
LaViLa lavila 35.1 26.6 30.8 33.7 30.4 32.0 30.9 35.1 93.6 59.1
AVION zhao2023avion 37.1 28.7 32.9 34.4 31.0 32.7 38.6 42.3 94.4 62.1
HelpingHands zhang2023helping 35.6 26.8 31.2 34.7 31.7 33.2 29.4 35.3 93.2 58.8
HENASY phan2024henasy 35.5 27.1 31.3 34.6 31.7 33.2 29.6 35.9 94.1 61.3
EgoDTM (ours) 37.9 29.1 33.5 34.8 31.9 33.4 40.2 43.2 94.6 63.6
@@ 4 Experiments
To evaluate our EgoDTM, we conduct experiments from three perspectives: short video understanding (video-text retrieval and action recognition), 3D space comprehension (depth estimation and robot manipulation), and long video understanding (natural language query and moment query). These evaluations cover seven benchmarks across five datasets. In the following sections, we detail the experimental setups (Section 4.1), main results (Section 4.2), ablation studies (Section 4.3), and further analyses (Section 4.4).
@@ 4.1 Benchmarks and Settings
Zero-shot video-text retrieval (ZS-VTR). To assess video-text alignment, we conduct zero-shot retrieval evaluations on text-to-video multi-choice retrieval on Ego4D grauman2022ego4d (EgoMCQ) and multi-instance video-text retrieval on Epic-Kitchens-100 epic-kitchens-100 (EK100MIR). Following lin2022egovlp; epic-kitchens-100, we use mean Average Precision (mAP) and the normalized Discounted Cumulative Gain (nDCG) as metrics for EK100MIR and accuracy for EgoMCQ.
Zero-shot action recognition (ZS-AR). Action recognition is conducted in a video-to-text retrieval manner. We experiment on the EGTEA egtea and Epic-Kitchens-100 epic-kitchens-100 (EK100CLS). The task on EGTEA requires models to recognize 106 classes of cooking activities, and the EK100CLS includes evaluation on 97 verbs and 300 nouns in kitchens. The metrics include mean accuracy, top-1 and top-5 accuracy across test splits.
Depth estimation (DE). To assess the model’s 3D awareness, we conduct depth estimation by finetuning our model on the H2O Kwon2021h2o dataset. Following cvpr-2024probing3D, we add a prediction head composed of a linear layer and an upsampling convolution layer on top of the frozen video encoder. Similar to cvpr-2024probing3D; eigen2014depth, the metrics are threshold accuracy ($\delta_{i}$): percentage of pixels $\tilde{D}_{j}$ in ground-truth depth maps that satisfy $\text{max}(\frac{D_{{\rm gt}j}}{\tilde{D}_{j}},\frac{\tilde{D}_{j}}{D_{{\rm gt}\ j}})=\delta_{i}<thr^{i}$ for $thr=1.25,i\in\{1,2,3\}$, and RMSE: $\|\tilde{D}-{D}_{\rm gt}\|_{2}$. The scale-invariant metrics are shifted and normalized from scale-aware metrics.
Table 2: Comparisons of depth estimation task on H2O dataset.
@@
ConvNext liu2022convnext 0.721 0.965 0.991 0.644 0.727 0.969 0.996 0.593
CLIP clip 0.795 0.966 0.988 0.624 0.811 0.976 0.994 0.559
EgoVLP lin2022egovlp 0.778 0.954 0.989 0.610 0.853 0.977 0.996 0.497
LaViLa lavila 0.801 0.954 0.987 0.598 0.811 0.964 0.993 0.552
AVION zhao2023avion 0.786 0.960 0.991 0.606 0.812 0.969 0.996 0.543
@@
Robot Manipulation (RM). The robot must learn to accurately perceive and interact in 3D spaces to accomplish daily tasks. Following previous works nair2022r3m; zeng2024mpi, we evaluate the visual representations as frozen perception modules for downstream policy learning within the Franka Kitchens simulation environment gupta2019frankakitchen. Five tasks are adopted: turn knob (TK), open door (OD), flip switch (FS), open microwave (OM), and slide door (SD). We use success rate as the metric.
Natural language queries (NLQ). The NLQ task localizes time intervals in a long video given a language query. We experiment on the EgoNLQ task on Ego4D grauman2022ego4d. To fairly evaluate different pretrained visual representations, we extract the visual features by video encoders and text features by the same pretrained BERT devlin2019bert, then train a VSLNet zhang2020vslnet to solve this task. The evaluation metrics are “Rn@m”, where $n\in\{1,5\}$ and $m\in\{0.3,0.5\}$, presenting the percentage of at least one of the top-$n$ predicted intervals having IoU greater than $m$.
Moment Queries (MQ). The MQ task is a video-only problem that aims to detect all temporal activity intervals in a long video given a specified activity category. We conduct experiments on the EgoMQ benchmark from Ego4D grauman2022ego4d. Following the setup in lin2022egovlp, we extract visual features using video encoders and then train a VSGN model zhao2021vsgn to perform this task. The evaluation metrics are “R@n, m”, consistent with those used in the NLQ task.
Pretraining details. The dual-encoders are initialized by the checkpoint pretrained on the original four million video-text pairs lin2022egovlp from Ego4D. EgoDTM is then trained for two epochs on 8*A800 GPUs, which requires approximately 10 hours and a learning rate of 3e-5. The depth-aware transformer decoder comprises six layers. The hidden dimension of the dual encoders is 768, while the 3D-aware decoder uses a dimension of 256 for efficient design. We use frames with 224p as input and 56p as output of the depth maps. Consequently, our 3D-aware decoder only has 9M parameters, and the batch size is set to 4096.
4.2 Main Results
Zero-shot video-text retrieval. In Table 1, EgoDTM outperforms models like LaViLa lavila and AVION zhao2023avion. For example, on EK100-MIR, EgoDTM outperforms the state-of-the-art AVION by +0.9% in mAP and +0.3% in nDCG, while also achieving a +1.5% improvement in intra accuracy on EgoMCQ. This indicates that captions generated based on visual cues from our data curation pipeline are more accurate and informative than captions generated from visual-conditioned LLMs or simply text rewriting. Besides, our model surpasses HelpingHands zhang2023helping and HENASY phan2024henasy, which rely on noisy HOI detection supervision with extra parameters in the visual encoder, indicating that depth modality can potentially offer larger merits for video-language models.
Zero-shot action recognition  Table 1 illustrates the superior performance of EgoDTM on EGTEA. While previous works have demonstrated that depth modality enhances traditional action recognition tasks zhang2016rgbd; liu2019ntu, their applicability has been largely constrained to specific scenarios and closed-set datasets. Our results extend this understanding, showing that video-language models can effectively leverage depth modality to learn more generalizable video representations.
Depth estimation. We evaluate the ability of visual-language models to infer depth from images in Table 2. Our model surpasses previous video-language models across most metrics by a large margin. In particular, our method improves the scale-aware RMSE by approximately 9.8% over the second-best result achieved by LaViLa. Interestingly, EgoVLP performs better in scale-invariant $\delta_{1}$ metric, suggesting that it latently captures finer-grained pixel-level details. We hypothesize that this advantage stems from EgoVLP’s scene-aware negative sampling strategy, which samples video-text pairs captured from the same environments into the same batch, thereby learning relevant details.
[…]
@@ 4.2 Main Results
R3M nair2022r3m 53.3% 50.7% 59.3% 86.3% 97.7% 69.4%
MPI zeng2024mpi 83.3% 54% 44.5% 93.5% 100% 75%
ResNet he2016resnet 28% 18% 26.7% 50% 75.5% 39.7%
CLIP clip 26.3% 13% 24.7% 41.7% 86.3% 38.4%
LaViLa lavila 48% 26% 22.5% 69% 94.5% 52%
@@
Robot manipulation. As shown in Table 9, EgoDTM consistently outperforms pretrained visual-language models such as CLIP clip and LaViLa lavila by +20.2% and +6.6%, respectively, demonstrating stronger spatial perception in visual representations. Additionally, EgoDTM performs competitively with specialized robot learning methods on certain tasks, such as “turn knob” but underperforms on others, like “open microwave”. A possible reason is that our model, pretained on depth and text, may overfit to real-worold scenarios, whereas methods like R3M nair2022r3m leverage self-supervised pretraining, which could provide better generalization to diverse manipulation tasks.
Moment queries. As shown in Table 8, EgoDTM achieves competitive performance among all methods. Specifically, it surpasses AVION by +2.78 on R5@0.3 and +2.85 on R5@0.5, and slightly improves over LaViLa by +1.4 on R5@0.3 and +1.56 on R5@0.5. These results demonstrate that depth-aware representation learning in EgoDTM effectively enhances spatial reasoning and temporal localization in long egocentric videos, leading to more accurate moment predictions.
Table 10: Zero-shot ablation studies. Models are pretrained on 4M video-text pairs from Ego4D and evaluated with zero-shot video-text retrieval on Epic-Kitchens-100-MIR epic-kitchens-100 and EgoMCQ lin2022egovlp. Unless specified, the default setting includes: ViT-B/16 backbone, batch size of 2048, random substitution of enriched texts, both $\mathcal{L}_{\rm depth}$ and $\mathcal{L}_{\rm augvtc}$ as losses, 8 depth queries, and 56p depth maps.
EK100MIR EgoMCQ
@@
(d) Model size. Large model brings more performance gain on intra accuracy of EgoMCQ.
EK100MIR EgoMCQ
@@ 4.3 Ablation Study
Table 11: Ablation studies on downstream tasks.
Metrics EK100MIR EgoMCQ EK100CLS EgoNLQ EgoMQ DE
mAP / nDCG $\uparrow$ inter / intra acc $\uparrow$ top-1 / top-5 acc $\uparrow$ mIoU $\uparrow$ mAP $\uparrow$ scale-aware RMSE / scale-invariant RMSE $\downarrow$
$\mathcal{L}_{\text{vtc}}$ 29.7 / 30.7 94.2 / 60.2 12.847 / 30.037 6.14 6.97 0.572 / 0.495
$\mathcal{L}_{\text{vtc}}+\mathcal{L}_{\text{depth}}$ 31.3 / 31.2 94.2 / 62.6 15.412 / 32.995 5.98 7.52 0.5637 / 0.464
$\mathcal{L}_{\text{augvtc}}$ 31.3 / 32.2 94 / 61.6 16.508 / 32.851 6.53 6.14 0.550 / 0.489
$\mathcal{L}_{\text{augvtc}}+\mathcal{L}_{\text{depth}}$ 33.1 / 33.1 94.6 / 62.6 15.898 / 33.895 6.17 8.87 0.539 / 0.481
EgoDTM’s components. Combining video-text matching with depth estimation enhances each learning objective. EgoDTM leverages depth maps to capture object relations, while textual descriptions enriched with shape and movement details can benefit more from depth information. As shown in the table Table 11, progressively adding our modules leads to consistent improvements across the majority of tasks, and our model outperforms all baselines in the main experiments. One potential limitation is that depth pretraining may negatively impact the performance on EgoNLQ to some extent. Nevertheless, our proposed AugVTC module is able to mitigate this effect and yields improvements that surpass those achieved by Avion, LaViLa, and the EgoVLP encoder as shown in Table4 of our main paper. Besides, comparing row 2 and row 4 when adding $\mathcal{L}_{\text{augvtc}}$, we observe an increase in the scale-invariant RMSE and a decrease in the scale-aware RMSE. We hypothesize that this may result from the proxy task bias introduced by multi-task pretraining.
Table 12: Evaluation of our HOI mask generation pipeline on HOI segmentation benchmark VISOR darkhalil2022visor in kitchen environment.
Method
mAP
@(0.5:0.95)
map
@0.50
hand-AP
@(0.5:0.95)
object-AP
@(0.5:0.95)
Upperbound darkhalil2022visor 60.70 73.31 90.87 30.53
Our Pipeline 43.02 54.88 61.01 25.03